# TASTE: Text-Aligned Speech Tokenization and Embedding for Spoken Language Modeling

### v2

@@ 2.2.2 Modeling TASTE Embedding
Besides the token code sets, recent progress on latent modeling [17, 28] motivates us to conduct experiments on modeling our text-aligned speech embedding. Referencing MELLE [28], we employ a linear layer that predicts the mean vector $\mu_{i}$ and a log-magnitude variance vector log $\text{log }\sigma^{2}_{i}$, where $i$ indicates the $i$-th frame of the sequence. And the final predicted latent of frame $i$ is denoted as $e_{i}=\mu_{i}+\sigma_{i}\odot\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$. Following MELLE, the straight-through estimator is applied to allow gradients to back-propagate properly during training.
@@ 3 Experiment Setup
3.1 Model Configuration
For our TASTE speech tokenizer, we initialize our encoder from Whisper [35]. Specifically, we use whisper-large-v3 for our initialization. By doing so, we can reduce computational cost between obtaining the ASR transcription and extracting the TASTE tokenization with the TASTE encoder frozen during training. On the other hand, we use the S3 token from CosyVoice [7] as the target unit for speech reconstruction. Since their speech tokenization facilitates additional speaker embedding, we follow the same procedure to obtain one. Adding speaker embedding allows global speaker voice control, which is a reasonable and useful scenario for spoken language models. The unit-to-speech vocoder is comprised of a flow model [23, 27] and a HifiGAN. We use the published pre-trained ones from  [7], and they are not involved in our training. For the quantizer, we set the RVQ layer $R=4$, the codebook size $512$, and the codebook dimension to be $256$. For the spoken language modeling, we follow previous work [11, 33, 6, 22] and initialize our spoken language model from a text LLM. However, this introduces the vocabulary mismatch problem between the ASR and LLM. We resolve this issue by using word-level TASTE tokenization and embedding, which is detailed in Appendix A.2. Moreover, we conduct Low-Rank Adaptation (LoRA) for parameter-efficient fine-tuning of our TASLMs. We set the corresponding hyperparameters rank $r=64$ and $\alpha=128$.
3.2 Dataset
We use two datasets–Emilia and LibriTTS–as our training datasets. Emilia [12] is an in-the-wild dataset where the speech is web-scaled and the transcriptions are pseudo-labeled. We use only the English subset of this multi-lingual corpus, which is about 40,000 hours. LibriTTS [50] is a reading-style corpus based on LibriSpeech [34]. We use all the training splits in LibriTTS for training, which is approximately 600 hours of speech. In addition, the test-clean split in LibriSpeech is used for evaluation purposes for our TASTE tokenizer and TASLMs.
@@ 4 Result
We separate the evaluation into two phases: Section 4.1 shows the results regarding our TASTE tokenization; while Section 4.2 evaluates our TASLM across multiple aspects, including acoustic, semantic, and continuation. For clarity, the metrics are introduced within each section.
4.1 Results of TASTE Tokenization
4.1.1 Speech Reconstruction Evaluation
We first represent the speech reconstruction evaluation results. For comprehensive evaluation, we use different metrics, including the reference-free metrics for quality assessment, and the reference-based metrics for evaluating the similarity between the reconstructed and the original speech.
Quality Assessment
We use ASR-WER, UTMOS [38], and DNS-MOS [37] as our metrics for evaluating the speech quality. For ASR-WER, we use HuBERT-Large [14] as the ASR model to transcribe the speech, and then calculate the word-error rate (WER) on the transcription.  [^1] https://huggingface.co/facebook/hubert-large-ls960-ft UTMOS and DNS-MOS are both neural-based MOS predictors. While both evaluate the speech quality, the design purpose of DNS-MOS makes it more suitable for evaluation regarding the noise levels.
Similarity Assessment
@@ 4.1.1 Speech Reconstruction Evaluation
For similarity, we measure ViSQOL, duration consistency (Drtn. Con.), speaker similarity (Spkr. Sim.) and the MUSHRA test as human evaluation. We use ViSQOL [4]is a production-ready tool that predicts speech quality via spectro-temporal image similarity comparisons. For the duration consistency, we first get the word-level alignment of the transcriptions of the original and the reconstructed speech using Montreal Forced Aligner [26]; then we calculate if the duration between each of the same words is matched under a preset tolerance window, which is set to 50 milliseconds. For MUSHRA human listening test, we reference the original protocal [40] to instruct evaluators to rate the similarity and quality on a scale of 1 to 100 with reference given.
@@
Table 1: The speech tokenization evaluation results on the test-clean split of LibriTTS. The evaluation is separated into the Quality and the Similarity assessments, as introduced in Section 4.1.1. We use gray text to indicate the worst-performing methods in each metric.
@@
Method Bitrate Quality Similarity WER $\downarrow$ UTMOS DNSMOS ViSQOL Drtn. Con. Spkr. Sim. MUSHRA Ground Truth 256k 2.1% 4.09 3.84 - - - 76.6 Encodec [5] 1500 5.1% 1.58 3.26 3.46 0.94 0.63 - 3000 2.6% 2.35 3.48 3.81 0.96 0.78 25.6 500 5.2% 1.27 2.99 2.80 0.94 0.35 - SpeechTokenizer [51] 2000 3.0% 3.56 3.60 3.65 0.97 0.80 53.9 4000 2.5% 3.90 3.76 4.03 0.98 0.92 - Mimi [6] 1000 3.1% 3.60 3.60 3.62 0.96 0.82 67.6 S3 token (topline) [7] 600 3.0% 4.18 3.90 3.30 0.96 0.82 70.2 Text-only (baseline) $\sim$50 5.9% 4.31 4.11 2.44 0.57 0.78 42.6 TASTE (ours) $\sim$150 4.4% 4.29 4.10 3.05 0.91 0.80 68.3
Speech Reconstruction Results
The evaluation results of our speech reconstruction on LibriSpeech are shown in Table 1. We highlight that our TASTE speech tokenization has the lowest bitrate among all the other speech tokenization methods. Note that since the speech tokenization is dynamic in frequency, we estimate our bitrate based on calculating the overall token count and the duration over the testing set. Despite the low bitrate, we generally attain much better performance comparing with the worst performing methods (gray text in the table) across each metric. Moreover, on the quality assessment, our MOS prediction scores are the second highest and even surpasses the ground truth, showcasing the reconstructed speech is of high quality. Next, we focus on the results of the similarity assessment. As for the duration consistency, we score the second-worst performance comparing with the other methods. We attribute this to the fact that our tokenization compress the sequence in a very dynamic way. Despite that, we still outperform the text-only method with a large margin, perform close towards other speech tokenization methods which all have a fixed down-sampling rate. Lastly, our method attains the second-highest MUSHRA score (excluding the ground-truth anchor). This highlights TASTE’s effectiveness: even without reproducing every microscopic detail, it still yields perceptually high-quality speech in human listening tests. Overall, TASTE carries rich paralinguistic information, facilitating high-end speech reconstruction under an extremely low bitrate.
@@ 4.1.2 TASTE for Text-Aligned Speech Editing
After comprehensive speech-reconstruction experiments, we show that TASTE can also perform text-aligned speech editing. Suppose we have two utterances with the same transcript but different paralinguistic characteristics. By exchanging their TASTE token sequences word by word, we ask whether the associated paralinguistic traits are transferred as well. To make the effect easy to see, we choose utterances that differ mainly in speaking rate and focus on duration changes. The overall text-aligned editing procedure is describe as follows: 1) Extract the TASTE tokens $\hat{\bm{z}}^{\text{orig}}$ for each source utterance 2) Swap the tokens at the desired text positions, resulting in edited TASTE tokens $\hat{\bm{z}}^{\text{edit}}$. 3) Decode the edited token sequence $\hat{\bm{z}}^{\text{edit}}$ back to speech. In Figure 3, we present the alignments obtained from MFA [26] of the original speech and the speech after editing, and compare them horizontally. As shown in the figure, words whose tokens were swapped exhibit clear duration shifts, while the untouched words keep their original timing—evidence that TASTE enables precise, text-aligned manipulation. Additional examples that target other paralinguistic dimensions are provided on our demo page.
4.2 Evaluating Text-Aligned Spoken Language Modeling
To provide a comprehensive evaluation of our text-aligned spoken language modeling (TASLM), we first compare our pre-trained SLM with other methods through speech continuation and likelihood-based benchmarks in Section 4.2.1. Then, to further investigate the understanding capabilities, we conduct evaluation with spoken question answering in Section 4.2.2.
Table 2: Pretrained SLM speech continuation and likelihood-based next-speech selection results. The superscripts at the bottom of the table indicate the base models used by each SLM, indicated by superscripts. Cascade models refer to the pipeline with ASR (whisper-large-v3), text continuation by LMs, and TTS (CosyVoice). This comparison evaluates SLMs and cascade models in continuation evaluation. As shown in the table, TASLM tends to preserve the semantic capabilities of LMs.

### v3

@@ 3.2.2 Modeling TASTE Embedding
Besides the token code sets, recent progress on latent modeling (Kim et al., 2024; Meng et al., 2024; Sun et al., 2024; Fan et al., 2025) motivates us to conduct experiments on modeling our text-aligned speech embedding. Referencing MELLE (Meng et al., 2024), we employ a linear layer that predicts the mean vector $\mu_{i}$ and a log-magnitude variance vector $\text{log }\sigma^{2}_{i}$, where $i$ indicates the $i$-th frame of the sequence. And the final predicted latent of frame $i$ is denoted as $e_{i}=\mu_{i}+\sigma_{i}\odot\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$. Following MELLE, the straight-through estimator is applied to allow gradients to back-propagate properly during training.
@@ 4 Experiment Setup
Model Configuration
For our TASTE speech tokenizer, we initialize our encoder from Whisper (Radford et al., 2023). By doing so, we can reduce computational cost between obtaining the ASR transcription and extracting the TASTE tokenization with the TASTE encoder frozen during training. On the other hand, we use the S3 token from Du et al. (2024a) as the target unit for speech reconstruction. Since their speech tokenization facilitates additional speaker embedding, we follow the same procedure to obtain one. Adding speaker embedding allows global speaker voice control, which is a reasonable and useful scenario for spoken language models. The unit-to-speech vocoder is comprised of a flow model (Lipman et al., 2022; Mehta et al., 2022) and a HifiGAN. We use the published pre-trained ones from  Du et al. (2024a), and they are not involved in our training. For the quantizer, we set the RVQ layer $R=4$, the codebook size $512$, and the codebook dimension to be $256$. For the spoken language modeling, we follow previous work (Hassid et al., 2023) and initialize our spoken language model from a text LLM. However, this introduces the vocabulary mismatch problem between the ASR and LLM. We resolve this issue by using word-level TASTE tokenization and embedding, which is detailed in Appendix A.4. Moreover, we conduct LoRA fine-tuning of our TASLMs, with hyperparameters rank $r=64$ and $\alpha=128$.
Dataset
We use two datasets–Emilia and LibriTTS–as our training datasets. Emilia (He et al., 2024) is an in-the-wild dataset where the speech is web-scaled and the transcriptions are pseudo-labeled. We use only the English subset of this multi-lingual corpus, which is about 40,000 hours. LibriTTS (Zen et al., 2019) is a reading-style corpus based on LibriSpeech (Panayotov et al., 2015). We use all the training splits in LibriTTS for training, which is approximately 600 hours of speech. In addition, the test-clean split in LibriSpeech is used for evaluation purposes for our TASTE tokenizer and TASLMs.
@@ 5 Result
We separate our experimental results into two parts. Section 5.1 discusses how TASTE strikes a good reconstruction quality while enables effective joint spoken language modeling; while Seciton 5.2 presents the additional results and ablation study of our joint tokenization and text-aligned SLM.
5.1 Main Results
To demonstrate the benefits of our joint tokenization, we first evaluate the performance of TASTE on speech reconstruction; then introduce how it allows effective spoken language modeling. For simplicity, the evaluation metrics are introduced within each section.
5.1.1 TASTE for Speech Reconstruction
Evaluation
@@ 5.1.1 TASTE for Speech Reconstruction
We evaluate our joint tokenization on two aspects: Quality and Similarity. For Quality assessment, we use ASR-WER, UTMOS (Saeki et al., 2022), and DNS-MOS (Reddy et al., 2021) as our metrics for evaluation. In ASR-WER, we use HuBERT-Large (Hsu et al., 2021) as the ASR model to transcribe the reconstructed speech, and then calculate the word-error rate (WER) on the transcription.  [^1] https://huggingface.co/facebook/hubert-large-ls960-ft UTMOS and DNS-MOS are both neural-based MOS predictors. While both evaluate the speech quality, the design purpose of DNS-MOS makes it more suitable for evaluation regarding the noise levels. For Similarity assessment, we measure ViSQOL, duration consistency (Drtn. Con.), speaker similarity (Spkr. Sim.), and the MUSHRA human listening test score. ViSQOL (Chinen et al., 2020)is a production-ready tool that predicts speech quality via spectro-temporal image similarity comparisons. For the duration consistency, we first get the word-level alignment of the transcriptions of the original and the reconstructed speech using Montreal Forced Aligner (McAuliffe et al., 2017); then we calculate if the duration between each of the same words is matched under a preset tolerance window, which is set to 50 milliseconds. In the MUSHRA human listening test, we follow the original protocal (Series, 2014) to instruct evaluators to rate similarity and quality on a scale of 1 to 100 with reference given.
@@
Table 1: The speech tokenization evaluation results on the test-clean split of LibriSpeech. The evaluation is separated into the Quality and the Similarity assessments, as introduced in Section 5.1.1. We use gray text to indicate the worst-performing methods in each metric. Freq. indicates the number of tokens per second. All reported results already account for the effect of ASR errors whenever textual transcriptions are involved (Text-only and TASTE).
@@
Method Freq. Bitrate Quality Similarity WER $\downarrow$ UTMOS DNSMOS ViSQOL Drtn. Con. Spkr. Sim. MUSHRA Ground Truth 16k 256k 2.1% 4.09 3.84 - - - 76.6 Encodecα 75 1500 5.1% 1.58 3.26 3.46 0.94 0.63 - 75 3000 2.6% 2.35 3.48 3.81 0.96 0.78 25.6 50 500 5.2% 1.27 2.99 2.80 0.94 0.35 - SpeechTokenizerβ 50 2000 3.0% 3.56 3.60 3.65 0.97 0.80 53.9 50 4000 2.5% 3.90 3.76 4.03 0.98 0.92 - Mimiγ 12.5 1000 3.1% 3.60 3.60 3.62 0.96 0.82 67.6 S3 tokenθ (topline) 25 600 3.0% 4.18 3.90 3.30 0.96 0.82 70.2 Text-only (baseline) $\sim$3 $\sim$50 5.9% 4.31 4.11 2.44 0.57 0.78 42.6 TASTE (ours) $\sim$3 $\sim$150 4.4% 4.29 4.10 3.05 0.91 0.80 68.3 α Défossez et al. (2023), β Zhang et al. (2024), γ Défossez et al. (2024), θ Du et al. (2024a)
Results Analysis
Table 1 reports the results of speech reconstruction on LibriSpeech. To better understand the effectiveness of TASTE, we highlight three main observations. 1) Since our tokens are text-aligned, TASTE operates at the lowest frequency and bitrate among all tokenization methods. We estimate these dynamic values by counting the total number of tokens and accumulating the duration over the testing set. 2) Despite this extremely low bitrate, TASTE achieves on-par or even superior performance to higher-bitrate methods in the quality assessment. In particular, TASTE yields lower ASR-WER than the text-only baseline, which we attribute to speech tokens carrying paralinguistic information that improves the naturalness of reconstructed speech. 3) In terms of similarity, TASTE performs comparably to high-bitrate, fixed down-sampling methods across multiple metrics. The inferior results on ViSQOL can be partly attributed to our use of a flow-based vocoder, as both TASTE and the S3 token topline exhibit weaker ViSQOL performance—a phenomenon also observed in Liu et al. (2025). This degradation on ViSQOL is not reflected in the MUSHRA listening test, where TASTE attains competitive perceptual quality and similarity from a human perspective. In general, TASTE significantly outperforms the text-only baseline, confirming that it carries sufficient paralinguistic information to allow high-quality speech reconstruction.
5.1.2 TASTE for Spoken Language Modeling
Table 2: Pretrained SLM speech continuation and likelihood-based next-speech selection results. The superscripts at the bottom of the table indicate the base models used by each SLM, indicated by superscripts. Cascade models refer to the pipeline with ASR (Radford et al., 2023), text continuation by LMs (Touvron et al., 2023), and TTS (Du et al., 2024a). This allow us to evaluate SLMs with cascade models in continuation perspective.
Method Finetuned / base Continuation Likelihood parameters GPT-4o UTMOS Human SALMON StoryCloze Overall Cascade Cascade (LLaMA3.2-1Bα) - 3.15 4.25 4.00 - - - Cascade (LLaMA2-7Bβ) - 3.43 4.25 3.98 - - - Spoken LMs TWIST 1.3B (Hassid et al., 2023) 1.3B / 1.3Bθ 1.48 3.25 1.95 62.5 61.5 62.0 TWIST 7B (Hassid et al., 2023)    7B /    7Bγ 1.44 3.27 2.04 63.4 64.7 64.1 Spirit LM (Nguyen et al., 2025)    7B /    7Bβ 2.79 3.41 2.38 59.1 72.0 65.6 Spirit LM Expr. (Nguyen et al., 2025)    7B /    7Bβ 1.90 3.40 2.41 69.0 66.2 67.6 Baseline (S3 token) 45M / 1.3Bα 1.37 4.04 2.84 50.2 58.7 54.5 TASLM 1B (token) 45M / 1.3Bα 3.08 4.07 3.93 60.8 76.5 68.7 TASLM 1B (embed.) 45M / 1.3Bα 3.16 4.22 4.16 57.7 76.7 67.2 Base models: αLLaMA3.2-1B, βLLaMA2-7B, γLLaMA-7B, θOPT-1.3B
TASTE is designed specifically to enable effective joint spoken language modeling (SLM). To examine its effectiveness, we train pretrained SLMs on top of TASTE following the methodology in Section 3.2. In line with prior work (Nguyen et al., 2025; Lin et al., 2024), we evaluate these models from two perspectives: speech continuation evaluation and likelihood-based evaluation.
Speech Continuation Evaluation
First, each pretrained SLM is conditioned on 3-second speech segments from LibriSpeech test-clean to generate speech continuations under their own decoding schemes, following Hassid et al. (2023); Lin et al. (2024). The generated continuations are then evaluated along two main aspects: semantic coherence and speech naturalness. For the semantic aspect, we transcribe the continuations using ASR and ask GPT-4o to assign MOS scores based on their coherence. For the speech naturalness aspect, we compute UTMOS as an objective score of speech quality. In addition, human evaluators provide an overall MOS score that jointly considers both coherence and naturalness. The detailed instructions given to GPT-4o and human evaluators are provided in Appendix A.3.2.
Likelihood-Based Evaluation
Following previous work (Hassid et al., 2023; Nguyen et al., 2025; Lin et al., 2024), we also evaluate the SLMs through likelihood-based benchmarks, where the accuracy score is based on whether the pretrained SLM chooses the correct continuation from the two given speech utterances based on its output likelihoods. We adopt two established benchmarks SALMON (Maimon et al., 2024) and spoken StoryCloze (Hassid et al., 2023; Mostafazadeh et al., 2016), which covers the acoustic aspect and the semantic aspect, respectively. Since both benchmarks contain multiple tasks, we report the average accuracy across these tasks within each benchmark for simplicity. The detailed results are in Appendix A.1.5 for the interested readers. We also report the mean of the SALMON and StoryCloze as an overall assessment for both aspects.
Results Analysis
The results of TASLM compared to other pre-trained SLMs are shown in Table 2, and three main advantages can be observed. 1) Compared to other pretrained SLMs, TASLM achieves substantially better performance on speech continuation across both human and machine evaluations, while also performing competitively on the likelihood-based benchmarks. Notably, this is achieved with only LoRA finetuning on a relatively small 1.3B base language model, illustrating the effectiveness of TASTE for joint modeling. 2) Compared to cascade models with the same base LM, our $\rm{TASLM}_{\text{emb}}$ achieves comparable scores on GPT-4o but higher human MOS. This indicates that its generated speech is more natural than cascade systems that rely solely on TTS during continuation. TASLM is the only SLM that not only maintains but even surpasses the performance of its corresponding text-based model, highlighting the importance of speech token modeling. 3) Directly using the S3 token for joint modeling following Xie and Wu (2024a) yields poor performance across all aspects, even though it surpasses TASTE in speech reconstruction. This shows that while reconstruction quality is critical, it is not the sole consideration in tokenization for spoken language modeling. Taken together, these results highlight the central contribution of TASTE: building a joint tokenization that facilitates more effective joint spoken language modeling.
5.2 Additional Results
@@ 5.2.1 TASTE for Text-Aligned Speech Editing
Beyond the main results presented above, we report several intriguing observations that further showcase the versatility of TASTE. The first is that TASTE naturally enables text-aligned speech editing, as illustrated in Figure 3. Suppose we have two utterances with the same transcript but different paralinguistic characteristics. By exchanging their TASTE token sequences word by word, we ask whether the associated paralinguistic traits are transferred as well. To make the effect clear, we select utterances that differ mainly in speaking rate and examine duration changes using MFA (McAuliffe et al., 2017). As illustrated in Figure 3, swapping tokens at specific word positions causes the corresponding words to exhibit clear duration shifts, while untouched words preserve their original timing—evidence that TASTE enables precise, text-aligned manipulation. This observation also echoes our design principle introduced in Section 1: a speech token should avoid redundantly encoding text content and instead concentrate on conveying paralinguistic information. Additional examples targeting other paralinguistic dimensions are provided on our demo page.
Table 3: Evaluation of spoken question answering. Performance across modalities is compared row-wise, where T is text and S denotes speech.
Method Mode Web Q. LLaMA-Q.
Mini-Omni 0.5B($\xrightarrow{}T$) T 21.3 39.0
Mini-Omni 0.5B T+S 4.5 11.6
Helium 7B (text) T 32.3 75.0
Moshi 7B T+S 26.6 62.3
LLaMA3.1-8B-Instruct T 60.4 71.7
Llama-Omni-8B T+S 35.5 67.3
LLaMA3.2-1B T 24.0 51.0
TASLM 1B (embed.) T+S 27.1 57.6
 We apply few-shot learning to facilitate question answering.
Table 4: Ablation study on the effects of each module in TASTE speech tokenizer. Enc. is encoder, Agg. is aggregator, and Quan. is quantizer. *: top-5 accuracy.
Modules Freq. S3 token Acc.*
Enc. 50Hz 0.98
Enc. + Agg. $\sim$3Hz 0.88
Enc. + Agg. + Quan. $\sim$3Hz 0.76
Enc. (last) 50Hz 0.84
Enc. + Agg. (last) $\sim$3Hz 0.78
Text-only $\sim$3Hz 0.65
5.2.2 TASLM for Spoken Question Answering
Next, we intriguingly find out that our TASLM exhibits spoken QA ability under few-shot scenario (Brown et al., 2020). We are the only pretrained SLM in Table 2 that exhibits this capability. As a result, we compare it against other instruction-finetuned joint SLMs in Table 4 to better understand the performance. We use two spoken question answering benchmarks, Web Questions (Berant et al., 2013) and LLaMA-Questions (Nachmani et al., 2024), following Défossez et al. (2024). We report the accuracy of answer containment. To more comprehensively assess the impact of adding speech, we also report the performance of each system’s underlying base text LLM. Notably, TASLM is the only approach that preserves its base text LLM’s performance. We attribute this to TASTE’s joint tokenization strategy. Specifically, we employ a straightforward one-to-one mapping between text and speech tokens, which enables simple and effective joint modeling.
5.2.3 Ablation Study on TASTE Speech Tokenizer
We run an ablation on TASTE speech tokenizer and use S3 token top-5 reconstruction accuracy as a proxy for information retention. Table 4 first covers the module-wise ablations of our encoder, aggregator, and quantizer. The aggregator sharply reduces token rate with only a small drop in accuracy. Adding the quantizer lowers accuracy further, but performance is still well above the text-only baseline. Secondly, we show that using only the last hidden state $\mathbf{h}^{(L)}$ performs worse than using the shallow hidden states $\mathbf{h}^{(l)}$ (as values for the aggregator), confirming our design choice.
6 Conclusion
In this work, we propose Text-Aligned Speech Tokenization and Embedding (TASTE), to facilitate joint text-speech spoken language modeling. By aggregating proper encoder representation through the specialized cross-attention mechanism and taking the ASR model as initialization, we make the speech tokenization text-aligned in an end-to-end manner with no explicit word alignment required. With our text-aligned speech tokenization and embedding, joint text-speech modeling becomes straightforward and effective. We conduct extensive experiments demonstrating the benefits of developing a joint tokenization tailored for spoken language modeling. We anticipate that these findings encourage further research on more effective joint tokenization for generative modeling.
Limitation
Several limitations of our current work suggest promising avenues for future development. First, while our pretrained spoken language model generates high-quality audio continuations, it lacks mechanisms for turn-taking and instruction following; developing a dialogue system is a practical next step. Second, TASTE has so far been evaluated on English; confirming its generalizability across other languages remains future work. Third, our tokenization method is tailored for joint SLMs, and its applicability to other generative tasks remains underexplored. Fourth, our pipeline currently focuses on single-speaker speech with lexical content and does not explicitly handle multi-speaker, overlapping, or non-lexical events (e.g., laughter, coughing). Future work could support these capabilities by incorporating target speech extraction (Zmolikova et al., 2023) and non-lexical event tags. Finally, system latency and streaming performance are yet to be optimized for real-time applications. Overall, none of these limitations is a fundamental barrier; rather, they are natural extensions and research targets that will further enhance the versatility of TASTE framework.
Ethics Statement
TASTE enables the efficient development of spoken language models. It lowers the barrier to building speech systems and improves the accessibility and convenience of human–computer interaction. At the same time, it raises security concerns: systems built with TASTE can more easily mimic a person’s voice and synthesize convincing personalized speech. Moreover, TASTE’s text-aligned speech editing makes voice manipulation straightforward. Overall, TASTE offers clear utility for beneficial applications, but responsible deployment—paired with consent, provenance, and anti-abuse safeguards—is essential to mitigate misuse risks. On the other hand, this study includes human evaluations in the form of subjective listening tests. Annotators were recruited from Amazon Mechanical Turk and compensated fairly according to the platform’s recommended rates. All participants provided informed consent, and no personal information has been collected. The audio material used in the evaluation does not contain sensitive content. The study adheres to the ethical standards commonly adopted in speech perception research.
Acknowledgement
This work is partially supported by the National Science and Technology Council (NSTC) in Taiwan under Grant NSTC 113-2634-F-002-008, and Liang-Hsuan Tseng is supported by the NSTC Graduate Research Fellowship (NSTC-GRF). We also extend our gratitude to the anonymous reviewers for their constructive comments, which can be found at https://openreview.net/forum?id=6STb8DauN1.
Appendix A Appendix
A.1 Supplementary Results
Table 5: The ablation study on how the ASR affects the performance of our TASTE tokenizer regarding speech reconstruction. GT: ground-truth transcription.
Method WER UTMOS DNS-MOS ViSQOL Drtn. Con. Spkr. Sim. TASTE (w/ ASR) 4.4% 4.29 4.10 3.05 0.91 0.80 TASTE (w/ GT) 4.6% 4.24 4.08 3.06 0.91 0.81
Table 6: The ablation study on how the ASR affects our SLM on spoken QA.
Methods Web-Q LLaMA-Q TASLM (w/ ASR) 27.1 57.6 TASLM (w/ GT) 28.0 57.7
Table 7: The ablation study on using a different ASR model regarding the SLM continuation semantic evaluation. Overall, we do not observe significant relative performance difference.
Evaluation Models TWIST 1.3B TWIST 7B Spirit LM Spirit LM Expr. S3 token TASLM (token) TASLM (embed.) Whisper + GPT-4o 1.48 1.44 2.79 1.90 1.37 3.08 3.16 nvidia-parakeet + GPT-4o 1.38 1.49 2.76 2.06 1.42 3.20 3.37
A.1.1 Ablation Study on the Effect of ASR
Because our tokenization, SLM, as well as the evaluation using GPT-4o all rely on an ASR system to extract text transcriptions, we conduct several ablation studies to assess the impact of ASR on performance. 1) In Table 6 and Table 6, we study how the ASR affects the performance of our TASTE tokenizer on speech reconstruction and TASLM on spoken question answering. Our results indicate that on both the tokenization and the SLM stages, the performance drop introduced by the ASR errors are almost negligible, primarily attributed to the robustness of recent ASR systems. Note that we do not use any ground-truth transcriptions in the previous experiments in the main text. 2) We study how substituting the ASR used to produce transcripts before GPT-4o’s semantic-coherence evaluation affects the reported scores. As shown in Table 7, we use another ASR model named nvidia-parakeet (Sekoyan et al., 2025), which employs an RNN-T (Graves, 2012; Xu et al., 2023) backbone. The results indicate that there is no significant relative performance difference between using Whisper and nvidia-parakeet ASR systems. TASLMs achieve much better results in both evaluation setups compared to the other pretrained SLMs.
A.1.2 Aggregator Cross Attention Visualization
To understand whether the aggregator has learned the text-speech aligned pattern, we visualize its cross attention map across all the layers and heads in Figure 4 and Figure 5.
Figure 4: The cross attention map of the last layer in our aggregator. As illustrated, a lot of heads clearly demonstrate the text-speech aligned behavior.
Figure 5: The cross attention map of the first layer in our aggregator. As illustrated, the behavior is quite different from the last layer in Figure 4. We observe some special heads: Head 8th showns clear alignment; while head 19th lights up at the silence part.
A.1.3 Discussion on the Selection of Shallow Hidden Layer
In Section 3.1.1, we propose using the shallow hidden representation from Whisper encoder as the key in our specialized cross attention, as it carries rich acoustic information. In practice, we select $l=6$. In Table 4, this allows us to achieve a near-optimal reconstruction accuracy on the target S3 unit. To further justify this selection, we follow the analysis methodology of Pasad et al. (2021) and compute the Canonical Correlation Analysis (CCA; Hotelling (1992)) between each Whisper encoder layer and the target S3 unit embeddings. The resulting similarity curve is shown in Figure 6. The correlations peak at the 4-th to 8-th layers, indicating that these shallow layers encode representations most aligned with the S3 targets—supporting our design choice of using a shallow hidden state. Importantly, Table 4 also shows that while the choice of the shallow layer affects the attainable upper bound, it does not impose a strict limitation on our method: even using the final encoder layer, whose correlation to S3 is much lower, still yields reconstruction quality better than the text-only baseline.
Figure 6: Canonical Correlation Analysis (CCA) between each Whisper encoder layer and the S3 target embeddings. Layers 4–8 exhibit the highest correlation with the S3 units, indicating that shallow representations contain the most target-relevant acoustic information.
Table 8: The tokenizer robustness ablation study. We apply four different level of noise, with signal-to-noise ratio (SNR) ranging from 20dB (almost clean) to 5dB (very noisy). Our method substantially achieves good reconstruction quality across all noise levels.
Method Bitrate SNR=20dB (clean) SNR=15dB SNR=10dB SNR=5dB (noisy) WER $\downarrow$ Sim. $\uparrow$ Rank WER $\downarrow$ Sim. $\uparrow$ Rank WER $\downarrow$ Sim. $\uparrow$ Rank WER $\downarrow$ Sim. $\uparrow$ Rank Ground Truth 256k 2.3% - - 2.5% - - 3.6% - - 9.2% - - 500 93.7% 0.193 13 97.7% 0.201 12 98.6% 0.205 12 98.3% 0.202 13 DACα 1000 36.0% 0.292 11 55.0% 0.276 11 80.2% 0.278 11 94.5% 0.283 10 12000 2.5% 0.937 1 2.9% 0.924 1 4.7% 0.914 1 12.3% 0.907 1 500 98.7% 0.295 12 104.6% 0.264 12 107.1% 0.264 12 105.8% 0.283 12 DM-Codecβ 1000 33.9% 0.479 9 53.0% 0.447 9 77.4% 0.435 9 97.1% 0.415 9 4000 3.9% 0.737 5 6.4% 0.689 6 14.4% 0.647 6 38.5% 0.595 6 500 15.2% 0.334 9 33.9% 0.324 9 64.2% 0.301 9 91.7% 0.277 10 SpeechTokenizerγ 2000 7.3% 0.773 8 16.3% 0.734 8 39.4% 0.682 8 73.8% 0.600 8 4000 4.4% 0.864 2 8.1% 0.822 4 21.0% 0.765 4 49.2% 0.684 5 BigCodecδ 1040 10.1% 0.829 7 17.8% 0.785 7 33.5% 0.718 7 63.0% 0.625 6 Mimiϵ 1000 5.1% 0.804 6 7.8% 0.772 5 14.7% 0.726 4 33.6% 0.673 4 S3 tokenζ 600 3.9% 0.860 2 5.2% 0.841 2 8.1% 0.815 2 16.7% 0.779 3 TASTE (ours) $\sim$150 4.8% 0.842 4 5.3% 0.830 3 6.9% 0.815 2 11.1% 0.792 1 α Kumar et al. (2023); β Ahasan et al. (2025); γ Zhang et al. (2024); δ Xin et al. (2024b); ϵ Défossez et al. (2024); ζ Du et al. (2024a)
A.1.4 Robustness Ablation on TASTE Tokenizer
To evaluate our tokenizer robustness, we have conducted controlled noise level experiments on speech reconstruction. Specifically, we add 4 different levels of white noise to the original waveform and then evaluate the reconstruction performance. The noise levels are defined by signal-to-noise ratio (SNR) ranging from 20dB (almost clean) to 5dB (very noisy). We report two metrics: ASR-WER as an indicator of reconstruction quality, and speaker similarity (Sim.) as a measure of reconstruction similarity. For ease of comparison, we also provide the overall rank of each tokenizer considering both metrics. The results in Table 8 show that our tokenizer remains stable across different noise levels, demonstrating strong robustness. Notably, TASTE achieves the best ASR-WER under the noisiest condition. This suggests that the underlying ASR system is not a limiting factor for the general applicability of TASTE; rather, it serves as an effective and reliable source of semantic information for the reconstruction.
A.1.5 Details on SALMON and StoryCloze
Our detailed results on SALMON and StoryCloze are reported in Table 9. The introductions of the two benchmarks—SALMON and StoryCloze—are described below.